20210223 LBTB Science OSCO + Bin 4

01:45 David having trouble with GCS
  • Not collimated yet
  • OSCO
  • OCAM not live in Center Pupil Closed loop
  • Wait for timeout. Interrupt Command button has no affect
  • AdSec went into Failure
  • Recover Fail us proceeding. Success
  • SetFlatAO
  • Wind at 2.3, 72 Elevation
  • Center Pupil has not returned
  • Change OCAM binning to 2
  • Still not returning from Center Pupils
  • Stop/Start AdSec Arbitrator. Not returned
  • Stop/Start AO Arbitrator. FLAO GUI Reset. Failure back to IIF
  • Stop/Start Masterdiagnostics to try to fix Closed Loop Not Live
  • Luci having trouble sending new preset. Dave says LUCI is waiting for Collimation and abort buttons is not available.
  • Dave reset something. Ok Now

02:15 New Preset.
  • GCS not find star. Check pointing
  • New preset.
  • Auto mode.
  • mag sent 13.9, bin 2, See 14.8. Bin 3 881 Hz.
  • GOPT On 0.13
  • First LUCI image no NCPA. Turned it on manually
  • All is good
night_feb23.png

bin3_nightfeb23.png

02:36 50" offset. Y position at -100. Sky frame so performance not important.

TN # 20210224_022405,022407,_025149,_025609

Estimated SR from the AO telemetry: 0.25
Seeing stable around 1arcsec. DIMM measurements.

03:00 modalplot. Loop stable

bin3_0300feb23_set2.png

03:14 Done with AO.
  • AuxLoop process did not die smile Yeah!
  • Re-applied the flat at 28 elevation when preset canceled
  • Is this safe, Guido?

ibution
idl_129502         |INF|      3063|2021-02-24 03:27:46.051365|             MAIN > Restore old oversampling...
idl_129502         |INF|      3064|2021-02-24 03:27:46.058636|             MAIN > Done.
idl_129502         |INF|      3065|2021-02-24 03:27:46.123963|             MAIN > Restore control current integrator action....
idl_129502         |INF|      3066|2021-02-24 03:27:46.165861|             MAIN > Done.
idl_129502         |INF|      3067|2021-02-24 03:27:46.167860|             MAIN > All Done. Succefully recovered from SafeSkipFrames.
idl_129502         |INF|      3068|2021-02-24 03:27:46.992601|             MAIN > Loading shape: /local/aomeas/adsec_calib/flat/Applicable/LUCIFER.sav
idl_129502         |INF|      3069|2021-02-24 03:27:48.272428|             MAIN > Selected flat: /local/aomeas/adsec_calib/flat/Applicable/LUCIFER.sav
idl_129502         |INF|      3070|2021-02-24 03:27:49.005398|             MAIN > The threshold has been moved to 0.2 as the total forces are included in the computation of the modal filtering.
idl_129502         |INF|      3071|2021-02-24 03:27:49.005455|             MAIN > The threshold has been additionally moved to 0.5 to try flattening better the external edge.
idl_129502         |INF|      3072|2021-02-24 03:27:49.005498|             MAIN > Delta command modal filtering for reducing peak force to 0.30 [N]
idl_129502         |INF|      3073|2021-02-24 03:27:49.043130|             MAIN > total current added in the force vector for the computation of the modal filtering
idl_129502         |INF|      3074|2021-02-24 03:27:49.053522|             MAIN > Maximum mode index used to apply flat -> # 246
idl_129502         |INF|      3075|2021-02-24 03:27:49.054416|             MAIN > Min/Max delta command to apply for flattening -> #  -6.37696e-06 /    6.41394e-06 [m]
idl_129502         |INF|      3076|2021-02-24 03:27:49.054601|             MAIN > Piston removed from commands
idl_129502         |INF|      3077|2021-02-24 03:27:49.257667|             MAIN > COMM  MAX [m] 6.86284e-05
idl_129502         |INF|      3078|2021-02-24 03:27:49.257709|             MAIN > FORCE MAX [N] 0.297268
AdSecArbitrator_41934|INF|       546|2021-02-24 03:27:49.718215|   COMMANDHANDLER > Command Load sha

-- DougMiller - 24 Feb 2021
Topic revision: r6 - 24 Feb 2021, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback